OpenStudy (anonymous):

An electric bulb is rated 220 V and 100 W. When it is operated on 110 V.Find the POWER.

terenzreignz (terenzreignz):

Well, power is directly proportional to voltage?

OpenStudy (anonymous):

yeah. I know it very well

terenzreignz (terenzreignz):

Well, I don't... I'm not even sure of that ^ :D Sorry~

OpenStudy (anonymous):

\[\huge P~ ~~\alpha ~~~~V ~and~ I\]

terenzreignz (terenzreignz):

Well, actually, \[\Large P \propto V\] if current is constant, but it may well not be... what might be constant is the resistance, in which case, given... \[\Large P = \frac{V^2}R\]\[\Large \color{blue}{P\propto V^2}\]instead...

OpenStudy (anonymous):

@terenzreignz why we use \(\large \frac{V^2}{R}\) ? Why don't we use only \(\large VI\)??

OpenStudy (anonymous):

any reason behind it ? @terenzreignz

terenzreignz (terenzreignz):

Because we know it's the same lightbulb, and therefore, the same resistance.

OpenStudy (anonymous):

but current is also the same in both the bulbs

OpenStudy (anonymous):

@Diwakar help me

OpenStudy (anonymous):

what is the confusion? @terenzreignz has showed how to solve it above. Power is proportional to the square of voltage. So, when u halve the voltage power becomes one-fourth.

OpenStudy (anonymous):

@Diwakar why can't we use \(\large VI\) formula ?

OpenStudy (anonymous):

tell me @Diwakar

OpenStudy (anonymous):

We are using this formula ,although implicitly. You see I=V/R. So, put this in your formula and see that you are getting P=V\(^{2}\)/R

OpenStudy (anonymous):

@Diwakar you use V^2/R formula. Why we can't use VI formula. It also finds the power. That's was my real problem

OpenStudy (anonymous):

these 2 formulas are same (at least in this problem). Look: P=VI =V *(V/R) =V^2/R

OpenStudy (anonymous):

the advantage of using this form is that R is constant no matter whether you are applying 220 or 110.

OpenStudy (anonymous):

when we use VI formula, the power is 50 W and when we use V^2/R formula, the power is 25 W. What will be correct answer?

OpenStudy (anonymous):

How are you getting 50W ? Let us see.First find resistance of bulb. As,P=V^2/R When you put V=220 and P=100. We get R=484 ohms. Now, let us find current when 110 V is applied I=V/R=110/484 P=VI=(110*110)/484 =\(\color{red}{25W}\)

OpenStudy (anonymous):

thank you :) for helping me.. @Diwakar

OpenStudy (anonymous):

no problem!

OpenStudy (anonymous):

@Diwakar sorry for disturbing you... so tell me that 50 W is wrong answer?

OpenStudy (anonymous):

25 is correct. 50 is wrong.

OpenStudy (anonymous):

thank you :)